# KVQuant: Towards 10 Million Context Length LLM Inference with KV Cache Quantization

### v3

@@ 3.6. Attention Sink-Aware Quantization
Prior work has demonstrated that after the first few layers in LLMs, the model tends to allocate a large attention score to the first token (xiao2023efficient, ). This occurs even when the initial token is not semantically important. This phenomenon happens because the model tends to use the inital token as a “sink”. (xiao2023efficient, ) demonstrates how keeping a few initial sink tokens allows for accurate performance for streaming applications. In our work, we demonstrate that due to the Attention Sink phenomenon, the model is disproportionately sensitive to quantization error in the first token. By keeping only the first token in fp16, we can attain significant perplexity benefits, particularly for 2-bit quantization. A similar observation has also been made in another concurrent work of (liu2024intactkv, ) in the context of weight quantization. Note that when retaining the first token in fp16, we account for this during the calibration process as well, meaning that we ignore the first token when deriving the nuqX datatype and when calibrating the scaling factors and zero points offline for the Keys. As demonstrated in Appendix I, this approach persistently yields performance benefits, particularly with lower bits and without dense-and-sparse quantization.
@@ 3.8. Kernel Implementation
In order to efficiently perform activation quantization on-the-fly, we leverage dedicated kernel implementations with our 4-bit quantization method for compressing vectors to reduced precision and extracting the sparse outliers, performing matrix-vector multiplications using the compressed vectors, and performing sparse matrix-dense vector multiplications using the sparse outliers. We store the quantized Key and Value matrices as 4-bit elements which are used as indices into lookup tables to recover the corresponding fp16 values. We store the sparse outlier matrices in either Compressed-Sparse Row (CSR) or Compressed-Sparse Column (CSC) format (depending on which aligns better with appending new Key and Value tokens). The kernels for the Key matrix-vector operations apply RoPE on-the-fly in order to support pre-RoPE quantization. More kernel implementation details are provided in Appendix Q.
@@ 4. Results
4.1. Main Results
4.1.1. Main Evaluation
We used the LLaMA-7B/13B/30B/65B, LLaMA-2-7B/13B/70B, and Mistral-7B models to evaluate our methodology by measuring perplexity on both Wikitext-2 and C4 (touvron2023llama, ; touvron2023llama2, ; jiang2023mistral, ); see Appendix K for details on our experimental setup. Table 2 shows the results for LLaMA models for the Wikitext-2 dataset. We compared our method with per-token quantization with and without grouping. The baseline configurations used by Atom and FlexGen are included for reference (zhao2023atom, ; sheng2023flexgen, ). We did not include results for KIVI (kivi, ) since they did not report perplexity values. We find that our method consistently outperforms baseline approaches by an especially large margin with 3-bit and 2-bit quantization. Once we incorporate outliers, we further push the performance of low-precision quantization, achieving 4-bit quantization with less than 0.02 perplexity degradation, 3-bit quantization with under 0.07 perplexity degradation, and 2-bit quantization with under 0.38 perplexity degradation on Wikitext-2, relative to the fp16 baseline, across all models (while attaining 3.7$\times$, 4.8$\times$, and 6.9$\times$ memory savings, respectively). Table 20 in Appendix M also provides zero-shot MMLU evaluation with 3-bit quantization, demonstrating how our method maintains performance on downstream tasks.
4.1.2. Long Context Length Evaluation
@@
We also evaluated the performance of our quantization method on passkey retrieval to assess the model’s ability to use its context. Passkey retrieval involves evaluating the model’s capacity to locate specific information in long texts (longchat2023, ), and this can be used to effectively measure the maximum distance over which a token can attend during the inference stage. We used the passkey evaluation framework from (zhu2023pose, ) (which is based on the methodology from (mohtashami2023landmark, )) to evaluate retrieval performance. Table 3 shows passkey retrieval results for the LLaMA-2-7B-32K and LLaMA-2-70B-32K models. These results indicate that with 4-bit and 3-bit quantization, KVQuant is able to maintain retrieval performance of the fp16 model. We observe some degradation with 2-bit quantization especially with LLaMA-2-70B, indicating potential room for improvement in 2-bit quantization techniques.
Table 3. Passkey retrieval results across different context lengths for the LLaMA-2-7B-32K model (uptrained for long sequence lengths using positional interpolation (chen2023extending, )) as well as the LLaMA-2-70B-32K LongLoRA model (chen2023longlora, ). The values reported are the success rate for retrieving the passkey, computed over 50 samples. 2-bit results are with per-matrix Q-Norm enabled.
Model Method 2K 4K 8K 16K 32K
LLaMA-2-7B-32K fp16 1 1 1 1 1
nuq4-1% 1 1 1 1 1
nuq3-1% 1 1 1 1 1
nuq2-1% 1 1 1 0.92 0.98
LLaMA-2-70B-32K fp16 1 1 1 1 1
nuq4-1% 1 0.98 1 1 1
nuq3-1% 1 0.98 1 1 1
nuq2-1% 0.82 0.8 0.82 0.8 0.76
4.1.3. Joint Weight and KV Cache Quantization
Table 4 shows results for our KV cache quantization method when the weights are also quantized using the methodology in SqueezeLLM (kim2023squeezellm, ). We observe minimal perplexity degradation when leveraging our KV cache quantization approach, even when weights are also quantized to reduced precision (achieving within 0.02 perplexity of 4-bit weight-only quantization when quantizing the KV cache using nuq4-1% for the LLaMA-7B and LLaMA-13B models). These results demonstrate how our method is compatible with existing weight-only quantization methods.
@@ 4.1.3. Joint Weight and KV Cache Quantization
fp16 6.13 5.45 16
w3-s45 nuq3 6.55 5.77 3.00
nuq3-1% 6.26 5.54 3.32-3.33
4.2. Performance Analysis
Table 5 shows kernel benchmarking results using a batch size of 1 for the 4-bit dense-and-sparse compression and matrix-vector kernel implementations. We show results across different sequence lengths to assess the performance of the kernels at different points during generation. We report latency benchmarked on an A6000 GPU and averaged over 1000 runs. The results show that for the Key and Value multiplications, we can achieve 1.1-1.2$\times$ and 1.2-1.4$\times$ latency savings, respectively, relative to the baseline. We have integrated these kernels into an end-to-end generation pipeline that is able to compress activations dynamically during inference, thereby achieving significant memory savings and allowing for either larger batch sizes or longer sequence lengths.
Table 5. Average latency (in microseconds) for the Key and Value nuq4-1% kernels, benchmarked on an A6000 GPU for the LLaMA-7B model across different sequence lengths ($l$). fp16 matrix-vector multiplication latencies are included for reference, and the fp16 Key multiplication time also includes the time to apply RoPE to the newly appended Key vector.  Section 3.8 and Appendix Q provide additional details for our kernel implementation, and Table 22 provides a detailed breakdown of kernel runtime.
@@ 4.2. Performance Analysis
Key fp16 Matvec 66.4 116.1 402.3
Key nuq4-1% 60.2 96.7 344.1
Value fp16 Matvec 56.0 104.2 389.3
Value nuq4-1% 40.8 85.8 293.4
4.3. Pushing the Context Length Limit
Table 6. Activation memory size (GB) for 128K, 1M, and 10M sequence length ($l$) for different LLaMA models. By compressing the KV cache to 2-bit precision, we can enable 1M context length inference with the LLaMA-7B model on a single A100-80GB GPU, and we can also enable 10M context length inference with the LLaMA-7B model on an 8-GPU system.
@@ 5. Conclusion
As context lengths in LLMs increase, the KV cache activations surface as the dominant contributor to memory consumption. Quantization is a promising approach to reduce the size of KV cache activations, but prior solutions failed to represent activations accurately in ultra-low precisions, such as sub-4-bit. In contrast, we achieve accurate ultra-low precision KV cache quantization. By quantizing Keys per-channel before applying RoPE, we are able to better match the outlier distribution and mitigate the impacts of RoPE on quantization (due to it mixing pairs of channels which may have different average magnitudes). We use non-uniform quantization to better allocate the small number of quantization signposts at low precision. We observe significant accuracy improvements when employing dense-and-sparse quantization, particularly when detecting outliers at the same granularity as we compute quantization scaling factors. Crucially, we demonstrate that we can perform accurate calibration offline for Keys, as well as efficient online scaling factor and outlier threshold computation for Values. By leveraging these methods, we are able to enable accurate low-precision activation quantization, achieving 4.8x compression (nuq3-1% outliers) with only 0.1 perplexity degradation across different LLaMA, LLaMA-2, and Mistral models. Additionally, we show that Q-Norm improves accuracy for 2-bit quantization by helping mitigate distribution shift. Our methodology therefore supports inferring the LLaMA-7B model with a context length of 10M on an 8-GPU serving system. Through our efficient kernel implementation, we are able to show improved latency relative to the fp16 baseline, demonstrating how our method allows for improved latency in addition to the memory savings.
@@ 6. Limitations
While our work enables accurate long-context length inference by reducing the memory requirements, there is significant work required for training long context length models with greater than 100K context length. This work is orthogonal to our efforts, which are constrained to efficient inference with long context length models. Additionally, our latency benchmarking results currently focus on memory-bandwidth bound generation rather than prompt processing (where we need to compress multiple Keys and Values at once). In future work, we plan to develop dedicated efficient kernels for block Key/Value compression in order to efficiently compress multiple Keys and Values at once. Finally, in the current end-to-end implementation, there are inefficiencies in how memory allocation is handled for updating the sparse matrix (where the data corresponding to the previous tokens have to be copied when concatenating them with the data from the new token). In future work, we plan to optimize this by doing blocked allocation to avoid overheads from reallocating memory.
@@ Appendix A RoPE Equation
The Query vectors computed at each iteration will have RoPE applied (to obtain $\tilde{Q}_{m}=R^{d}_{\theta,m}*Q_{m}$). When caching Key vectors, we therefore need to either cache $\tilde{K}_{n}=R^{d}_{\theta,n}*K_{n}$, or else we need to cache $K_{n}$ and apply $R^{d}_{\theta,n}$ on-the-fly during inference. In order to apply $R^{d}_{\theta,n}$ efficiently on-the-fly, we leverage the element-wise formulation of RoPE rather than the matrix-multiplication formulation from Equation 4. The element-wise formulation for $R^{d}_{\theta,n}x$ is as follows, where $\odot$ is the element-wise multiplication operator (note that the formulation that we use matches the implementation in Huggingface for LLaMA, and it is a different but equivalent formulation to the element-wise expression in [33]):
@@
Appendix B Key and Value Dynamic Range
Appendix C Derivation for Sensitivity Analysis
@@
In particular, the sensitivity measure is based on how much the loss output of the model is perturbed. To compute this we denote activations at a layer before quantization as $\bf A$, after quantization as $\bf A_{Q}$, quantization perturbation in activation as $\bf A-A_{Q}$, and the gradient of the Loss function w.r.t. activation as $J(A)=\frac{\partial{\mathcal{L}}}{\partial A}(A)$. By making the assumption that the quantization perturbation of different activations follow a Gaussian distribution with zero mean, we can show that the sensitivity of an activation value is proportional to $\mathcal{F}_{ii}\big{(}A-Q(A)\big{)}^{2}$ as was given in Eq. 2:
@@
Here note that we first assume a first order Taylor series expansion to approximate the perturbation to the loss, and then use the zero mean assumption of the quantization perturbation to derive the second line. To approximate the $\mathbb{E}_{\Delta{\bf A}}\left[\Delta A_{i}^{2}\right]$ we use empirical evaluation of the expectation by sampling multiple different inputs and empirically computing the resulting quantization perturbation which results in Eq. 2.
Appendix D Per-Channel Key Quantization Ablations
We report results in Table 7 demonstrating the perplexity for different KV cache compression ratios, showing that per-channel quantization for Keys and per-token quantization for Values outperforms the standard per-token quantization approach for both Keys and Values, yielding an improvement of 3.88 perplexity for the LLaMA-7B model at 3-bit precision. This demonstrates the benefits of per-channel Key quantization to mitigate the large outlier channels in Keys. Additionally, although there are per-channel outliers in Values, we observe that per-channel quantization for Values actually performs worse than per-token quantization. We hypothesize that this behavior is because per-channel Value quantization leads to greater error accumulation in particular output values (since the result of the attention scores multiplied by one channel of the Values will be localized to a single value in the output vector), which leads to greater quantization error at later model layers. Another concurrent work, KIVI [26], observes similar behavior for per-channel Value quantization, which they attribute to the fact that per-token Value quantization confines the error to each token. Assuming that the output is a weighted sum of only a few important tokens (as only a few attention scores are large), a perturbation in these tokens can lead to significant degradation.

### v4

@@ 3.5. Attention Sink-Aware Quantization
Prior work has demonstrated that after the first few layers in LLMs, the model tends to allocate a large attention score to the first token (xiao2023efficient, ). This occurs even when the initial token is not semantically important. This phenomenon happens because the model tends to use the inital token as a “sink”. In our work, we demonstrate that due to the Attention Sink phenomenon, the model is disproportionately sensitive to quantization error in the first token. By keeping only the first token in fp16, we can attain perplexity benefits, particularly for 2-bit quantization. A similar observation has also been made in another concurrent work  (liu2024intactkv, ). Note that when retaining the first token in fp16, we account for this during the calibration process as well, meaning that we ignore the first token when deriving the nuqX datatype and when calibrating the scaling factors and zero points offline for the Keys. As demonstrated in Appendix I, this approach persistently yields performance benefits, particularly with lower bit widths and without dense-and-sparse quantization.
@@ 3.7. Kernel Implementation
In order to efficiently perform activation quantization on-the-fly, we leverage dedicated kernel implementations with our 4-bit quantization method for compressing vectors to reduced precision and extracting the sparse outliers, performing matrix-vector multiplications using the compressed vectors, and performing sparse matrix-dense vector multiplications using the sparse outliers. We store the quantized Key and Value matrices as 4-bit elements which are used as indices into lookup tables to recover the corresponding fp16 values. We store the sparse outlier matrices in either Compressed-Sparse Row (CSR) or Compressed-Sparse Column (CSC) format (depending on which aligns better with appending new Key and Value tokens). The kernels for the Key matrix-vector operations apply RoPE on-the-fly in order to support pre-RoPE quantization. More kernel implementation details are provided in Appendix P.
@@ 4. Results
4.1. Main Evaluation
We used the LLaMA-7B/13B/30B/65B, Llama-2-7B/13B/70B, Llama-3-8B/70B, and Mistral-7B models to evaluate our methodology by measuring perplexity on both Wikitext-2 and C4 (touvron2023llama, ; touvron2023llama2, ; llama3modelcard, ; jiang2023mistral, ; wikitext2, ; c4, ); see Appendix K for details on our experimental setup, as well as our methodology for computing KV cache size estimates. Table 2 shows the results for LLaMA models for the Wikitext-2 dataset. We compared our method with per-token quantization with and without grouping. The baseline configurations used by Atom and FlexGen are included for reference (zhao2023atom, ; sheng2023flexgen, ). We find that our method consistently outperforms baseline approaches by an especially large margin with 3-bit and 2-bit quantization. Once we incorporate outliers, we further push the performance of low-precision quantization, achieving 4-bit quantization with less than 0.02 perplexity degradation, 3-bit quantization with under 0.1 perplexity degradation, and 2-bit quantization with under 0.5 perplexity degradation on Wikitext-2, relative to the fp16 baseline, across all models (while attaining 3.7$\times$, 4.8$\times$, and 6.9$\times$ memory savings, respectively).
4.2. Long Context Length Evaluation
@@
We also evaluated the performance of our quantization method on passkey retrieval to assess the model’s ability to use its context. Passkey retrieval involves evaluating the model’s capacity to locate specific information in long texts (longchat2023, ), and this can be used to effectively measure the maximum distance over which a token can attend during the inference stage. We used the passkey evaluation framework from (zhu2023pose, ) (which is based on the methodology from (mohtashami2023landmark, )) to evaluate retrieval performance. The passkey retrieval results are provided in Table 3, demonstrating how our method is able to maintain the retrieval performance for long context length models. We also include comparisons with the passkey retrieval when using KIVI for the LLaMA-2-7B-32K model (kivi, ), demonstrating that our approach can attain higher retrieval rate for the same compression level. Our improved performance on retrieval tasks can be attributed to our improved representation of all tokens equally (whereas KIVI preserves a local window of residual tokens in fp16, which better represents the tail part of the context but doesn’t provide as much benefit for tasks that need to utilize the full context window). Additionally, we include evaluation on Longbench in Appendix O, which contains a suite of long-context length evaluation benchmarks including QA tasks, summarization, and few-shot learning (bai2023longbench, ). We demonstrate that our 3-bit approach can attain minimal degradation relative to fp16 inference, and that our approach outperforms KIVI for a similar compression level. Note that for generative tasks, when processing the input prompt, the Key/Value matrix multiplications are computed using the fp16 Keys and Values, and then the Keys and Values are separately compressed into low precision.
Table 3. Passkey retrieval results across different context lengths for the LLaMA-2-7B-32K model (uptrained for long sequence lengths using positional interpolation (chen2023extending, )) as well as the Llama-2-70B-32K LongLoRA model (chen2023longlora, ). The values reported are the success rate for retrieving the passkey, computed over 50 samples. We also include comparisons with KIVI for reference, using the 2-bit configuration with group size of 32 and 128-element fp16 residual (kivi, ). Average bit widths are estimated for each approach assuming 32K context length. Note that the open-source code for running KIVI with LLaMA does not support grouped-query attention, so we did not include comparisons with KIVI for Llama-2-70B-32K.
Model Method 2K 4K 8K 16K 32K Bit Width
LLaMA-2-7B-32K fp16 1 1 1 1 1 16
KIVI-2 0.76 0.72 0.72 0.68 0.7 3.05
nuq4-1% 1 1 1 1 1 4.33
nuq3-1% 0.98 1 1 1 1 3.33
nuq2-1% 1 1 0.98 1 1 2.33
Llama-2-70B-32K fp16 1 1 1 1 1 16
nuq4-1% 1 1 1 1 1 4.35
nuq3-1% 1 1 1 1 1 3.35
nuq2-1% 0.98 0.98 0.96 1 0.74 2.35
4.3. Joint Weight and KV Cache Quantization
Table 4 provides results for our method when applied along with quantizing weights to low precision (where the weight quantization follows the approach in (kim2023squeezellm, )). We observe minimal perplexity degradation when leveraging our KV cache quantization approach, even when weights are also quantized to reduced precision (achieving within 0.02 and 0.1 perplexity of 4-bit and 3-bit weight-only quantization, respectively, when quantizing the KV cache using nuq4-1% for the LLaMA-7B and LLaMA-13B models). These results demonstrate how our method is compatible with existing weight-only quantization methods.
@@ 4.3. Joint Weight and KV Cache Quantization
w3-s45 fp16 6.13 5.45 16
nuq3-1% 6.23 5.52 3.32-3.33
4.4. Performance Analysis and Memory Savings
Table 5 shows kernel benchmarking results using a batch size of 1 for the 4-bit dense-and-sparse compression and matrix-vector kernel implementations. We show results across different sequence lengths to assess the performance of the kernels at different points during generation. We report latency benchmarked on an A6000 GPU. The results show that for the Key and Value multiplications, we can achieve 1.2-1.6$\times$ and 1.3-1.7$\times$ latency savings, respectively, relative to the baseline. We have integrated these kernels into an end-to-end generation pipeline that is able to compress activations dynamically during inference, thereby achieving significant memory savings and allowing for either larger batch sizes or longer sequence lengths.
Table 5. Average latency (in microseconds) for the Key and Value nuq4-1% kernels, benchmarked on an A6000 GPU for the LLaMA-2-7B-32K model across different sequence lengths ($l$). fp16 matrix-vector multiplication latencies are included for reference, and the fp16 Key multiplication time also includes the time to apply RoPE to the newly appended Key vector.  Section 3.7 and Appendix P provide additional details for our kernel implementation, Appendix P describes our benchmarking methodology, and Table 20 provides a detailed breakdown of kernel runtime on an A6000 GPU.
@@ 4.4. Performance Analysis and Memory Savings
Key fp16 Matvec 33.3 59.1 219.4
Key nuq4-1% 25.6 39.9 126.3
Value fp16 Matvec 26.0 50.2 203.7
Value nuq4-1% 22.1 37.9 124.5
Table 6. Activation memory size estimates (GB) for 128K, 1M, and 10M sequence length ($l$) for different LLaMA models. By compressing the KV cache to 2-bit precision, we can enable 1M context length inference with the LLaMA-7B model on a single A100-80GB GPU, and we can also enable 10M context length inference with the LLaMA-7B model on an 8-GPU system.
@@ 5. Conclusion
As context lengths in LLMs increase, the KV cache activations surface as the dominant contributor to memory consumption. Quantization is a promising approach to reduce the size of KV cache activations, but prior solutions failed to represent activations accurately in ultra-low precisions, such as sub-4-bit. In contrast, we achieve accurate ultra-low precision KV cache quantization. By quantizing Keys per-channel before applying RoPE, we are able to better match the outlier distribution and mitigate the impacts of RoPE on quantization (due to it mixing pairs of channels which may have different average magnitudes). We use non-uniform quantization to better allocate the small number of quantization signposts at low precision. We observe significant accuracy improvements when employing dense-and-sparse quantization, particularly when detecting outliers at the same granularity as we compute quantization scaling factors. Crucially, we demonstrate that we can perform accurate calibration offline for Keys, as well as efficient online scaling factor and outlier threshold computation for Values. By leveraging these methods, we are able to enable accurate low-precision activation quantization, achieving 4.8x compression (nuq3-1% outliers) with only 0.1 perplexity degradation across different LLaMA, Llama-2, Llama-3, and Mistral models. Our methodology therefore supports inferring the LLaMA-7B model with a context length of 10M on an 8-GPU serving system. Through our efficient kernel implementation, we are able to show improved latency relative to the fp16 baseline, demonstrating how our method allows for improved latency in addition to the memory savings.
@@ 6. Limitations
While our work enables accurate long-context length inference by reducing the memory requirements, there is significant work required for training long context length models with greater than 100K context length. This work is orthogonal to our efforts, which are constrained to efficient inference with long context length models. Additionally, our latency benchmarking results currently focus on memory-bandwidth bound generation rather than prompt processing (where we need to compress multiple Keys and Values at once). Finally, in the current end-to-end implementation, there are inefficiencies in how memory allocation is handled for updating the sparse matrix (where the data corresponding to the previous tokens have to be copied when concatenating them with the data from the new token). In future work, we plan to optimize this by doing blocked allocation to avoid overheads from reallocating memory.
@@ Appendix A RoPE Equation
The Query vectors computed at each iteration will have RoPE applied (to obtain $\tilde{Q}_{m}=R^{d}_{\theta,m}*Q_{m}$). When caching Key vectors, we therefore need to either cache $\tilde{K}_{n}=R^{d}_{\theta,n}*K_{n}$, or else we need to cache $K_{n}$ and apply $R^{d}_{\theta,n}$ on-the-fly during inference. In order to apply $R^{d}_{\theta,n}$ efficiently on-the-fly, we leverage the element-wise formulation of RoPE rather than the matrix-multiplication formulation from Equation 3. The element-wise formulation for $R^{d}_{\theta,n}x$ is as follows, where $\odot$ is the element-wise multiplication operator (note that the formulation that we use matches the implementation in the Transformers library for LLaMA [39], and it is a different but equivalent formulation to the element-wise expression in [34]):
@@
Appendix B Derivation for Sensitivity Analysis
@@
In particular, the sensitivity measure is based on how much the loss output of the model is perturbed. To compute this we denote activations at a layer before quantization as $\bf A$, after quantization as $\bf A_{Q}$, quantization perturbation in activation as $\bf A-A_{Q}$, and the gradient of the Loss function w.r.t. activation as $J(A)=\frac{\partial{\mathcal{L}}}{\partial A}(A)$. By making the assumption that the quantization perturbation of different activations follow a Gaussian distribution with zero mean, we can show that the sensitivity of an activation value is proportional to $\mathcal{F}_{ii}\big{(}A-Q(A)\big{)}^{2}$ as was given in Equation 2:
@@
Here note that we first assume a first order Taylor series expansion to approximate the perturbation to the loss, and then use the zero mean assumption of the quantization perturbation to derive the second line. To approximate the $\mathbb{E}_{\Delta{\bf A}}\left[\Delta A_{i}^{2}\right]$ we use empirical evaluation of the expectation by sampling multiple different inputs and empirically computing the resulting quantization perturbation which results in Equation 2.
Appendix C Derivation for Quantization Error
In our work, before applying the sensitivity-weighted K-means to derive quantization signposts, we normalize each element $A_{i}$ in the flattened activation $A$. This normalization for $A_{i}$ involves a shift by a zero-point $z_{i}$ followed by rescaling the quantization signposts by a scaling factor $s_{i}$, where $s_{i}$ and $z_{i}$ are the scaling factor and zeropoint corresponding to element $A_{i}$:
(5) $A_{i,norm}=\frac{A_{i}-z_{i}}{s_{i}},$
where $A_{i}$ and $A_{i,norm}$ are element $i$ from activation $A$ before and after normalization, respectively. We then quantize $A_{i,norm}$ to $Q(A_{i,norm})$ with quantization error $\Delta A_{i,norm}$. After we dequantized, we rescale each element by $s_{i}$ and add $z_{i}$ to get the recovered quantized activation value $Q(A_{i})$:
(6) $Q(A_{i})=s_{i}\>Q(A_{i,norm})+z_{i}.$
As such, if there is quantization error $\Delta A_{i,norm}$ in $A_{i,norm}$, this will be scaled by $s_{i}$ in terms of the error in $A_{i}$, i.e., $\Delta A_{i}=s_{i}\Delta A_{i,norm}$. For activation $A$ which is normalized to $A_{norm}$ (with corresponding scaling factors $s_{i}$ for each element $A_{i}$), minimizing the sensitivity-weighted quantization error as expressed in Equation 2 gives us the following expression, which we can minimize using the normalized activations across all $N$ elements from the samples in a calibration set:
(7) $Q(A)^{*}$ $\simeq\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\big{(}A_{i}-Q(A_{i})\big{)}^{2}$
(8) $=\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\Big{(}s_{i}^{2}\big{(}A_{i,norm}-Q(A_{i,norm})\big{)}^{2}\Big{)}$
Appendix D Key and Value Dynamic Range
Appendix E Per-Channel Key Quantization Ablations
As shown in Table 7, per-channel quantization for Keys and per-token quantization for Values outperforms the standard per-token quantization approach for both Keys and Values, yielding an improvement of 3.82 perplexity for the LLaMA-7B model at 3-bit precision. This demonstrates the benefits of per-channel Key quantization to mitigate the large outlier channels in Keys. Note that for all experiments using per-channel quantization, we use an fp16 zeropoint rather than a low-precision zeropoint that is rounded to the nearest integer value. We do this since for some of the Key channels, all of the elements are positive or all of the elements are negative, meaning that the zeropoint will fall outside of the range that is representable by a low-precision integer (and rounding it to the nearest low-precision value can degrade performance).
Additionally, we observe that per-channel quantization for Values actually performs worse than per-token quantization. We hypothesize that this behavior is because per-channel Value quantization leads to greater error accumulation in particular output values (since the result of the attention scores multiplied by one channel of the Values will be localized to a single value in the output vector), which leads to greater quantization error at later model layers. Another concurrent work, KIVI [25], observes similar behavior for per-channel Value quantization, which they attribute to the fact that per-token Value quantization confines the error to each token. Assuming that the output is a weighted sum of only a few important tokens (as only a few attention scores are large), a perturbation in these tokens can lead to significant degradation. Per-token Value quantization therefore ensures that the quantization of unimportant tokens does not adversely impact the important tokens.